OpenStudy (anonymous):

WHAT IS X,WHEN 6(X-6)=78/13?

OpenStudy (anonymous):

distribute: 6x-12= 78/13 add 12 from both sides: 6x= 234/13 divide 6 from both sides: x=39/13 simplify if u can: x=3